# CodeSense: a Real-World Benchmark and Dataset for Code Semantic Reasoning

### v1

@@ A.4 Detailed Description of Tasks
In this section, we will provide a detailed description of our tasks. A.4.1 provides a comprehensive description of the statement-based evaluation we performed on LLMs. We sampled five types of statements, i.e, Assignment, Arithmetic, Constant, Boolean, and Function Call, and prompted the models about the value after execution of each type of statement given the variable states before executing that statement. For the block prediction task A.4.2, we sampled statements from the start of the code snippet, given the input of the code, we prompted the model to predict the output at the end of the 1st statement, the 2nd statement, and the 3rd statement. For Branch prediction A.4.3, we promoted the model whether a specific branch will be taken or not of a code snippet, given the input of that code snippet.
@@
In the case of A.4.4, we sampled loop statements from the code snippets. For each loop, we first collected the number of iterations of that loop as ground truth, and queried the model about how many times the loop would be iterated. We also collected and prompted the model regarding the variable state inside the loop body after the n-th interaction. We have named this "In-Loop" prediction. Additionally, we sampled variables after the execution of the whole loop and queried the model regarding the variable value after the execution of the loop body ("Post-Loop" prediction). For input/output prediction A.4.5, we used the approach similar to  [8]. For output prediction, we give the entire code snippet and the input of the code snippet, and vice versa for input prediction.
A.4.1 Statement Prediction Task
@@
def xldate_from_date_tuple(date_tuple, datemode): year, month, day = date_tuple data_list = list(data_tuple) if datemode not in (0, 1): raise XLDateBadDatemode(datemode) if year == 0 and month == 0 and day == 0: return 0.00 c = 100 if not (1900 <= year <= $9$): raise XLDateBadTuple("Invalid_year:%r" % ((year, month, day),)) if not (1 <= month <= 12): raise XLDateBadTuple("Invalid_month:%r" % ((year, month, day),)) if day < 1 \ or (day > _days_in_month[month] and not (day == 29 and month == 2 and _leap(year))): raise XLDateBadTuple("Invalid_day:%r" % ((year, month, day),)) Yp = year + 4716 M = month if M <= 2: Yp = Yp - 1 Mp = M + 9 else: Mp = M - 3 jdn = ifd(1461 * Yp, 4) + ifd(979 * Mp + 16, 32) + \ day - 1364 - ifd(ifd(Yp + 184, 100) * 3, 4) xldays = jdn - _JDN_delta[datemode] if xldays <= 0: raise XLDateBadTuple("Invalid(year,month,day):%r" % ((year, month, day),)) if xldays < 61 and datemode == 0: raise XLDateAmbiguous("Before1900-03-01:%r" % ((year, month, day),)) return float(xldays) xldate_from_date_tuple(date_tuple=(1907, 7, 3), datemode=0) Assignment Prediction What will be the value of the statement year, month, day = date_tuple given {’date_tuple’: (1907, 7, 3)}? Arithmetic Prediction What will be the value of the statement Yp = year + 4716 given {’year’: 1907}? Constant Prediction What will be the value of the statement c = 0 given {’constant’: 0}? Boolean Prediction Will the statement if datemode not in (0, 1): be executed given {’datemode’: 0}? Function Call Prediction What will be the value of the statement data_list = list(data_tuple) given {’date_tuple’: (1907, 7, 3)}?
A.4.2 Block Prediction Task
@@
A.4.3 Branch Task
1.def xldate_from_date_tuple(date_tuple, datemode): 2. 3. year, month, day = date_tuple 4. 5. if datemode not in (0, 1): 6. raise XLDateBadDatemode(datemode) 7. 8. if year == 0 and month == 0 and day == 0: 9. return 0.00 10. 11. if not (1900 <= year <= $9$): 12. raise XLDateBadTuple("Invalid_year:%r" % ((year, month, day),)) 13. if not (1 <= month <= 12): 14. raise XLDateBadTuple("Invalid_month:%r" % ((year, month, day),)) 15. if day < 1 \ 16. or (day > _days_in_month[month] and not(day == 29 and month == 2 and _leap(year))): 17. raise XLDateBadTuple("Invalid_day:%r" % ((year, month, day),)) 18. 19. Yp = year + 4716 20. M = month 21. if M <= 2: 22. Yp = Yp - 1 23. Mp = M + 9 24. else: 25. Mp = M - 3 26. jdn = ifd(1461 * Yp, 4) + ifd(979 * Mp + 16, 32) + \ 27. day - 1364 - ifd(ifd(Yp + 184, 100) * 3, 4) 28. xldays = jdn - _JDN_delta[datemode] 29. if xldays <= 0: 30. raise XLDateBadTuple("Invalid(year,month,day):%r" % ((year, month, day),)) 31. if xldays < 61 and datemode == 0: 32. raise XLDateAmbiguous("Before1900-03-01:%r" % ((year, month, day),)) 33. return float(xldays) 34. 35.xldate_from_date_tuple((1907, 7, 3), 0) Brach Prediction Is line 12, raise XLDateBadTuple("Invalid year: %r" % ((year, month, day),)) executed when xldate_from_date_tuple((1907, 7, 3), 0) is called?
A.4.4 Loop Task
@@
A.4.5 Input-Output Task
@@
A.5 Concrete to Abstract Mapping
For the approximation of code semantics, we prompted the models to reason in abstract values, instead of reasoning about the concrete exact value. Table 6 shows the mapping from concrete value to abstract category, following the prior literature [9]. When defining these mappings, we carefully aligned the value ranges for each abstract category with the overall value distribution observed in our benchmark.
@@ A.5 Concrete to Abstract Mapping
$-10\leq v<0$ Negative Regular
$v<-10$ Negative Large
@@
$-1.0\leq v<0.0$ Negative Small
$-10.0\leq v<-1.0$ Negative Regular
$v<-10.0$ Negative Large
@@ A.6.1 RQ1 Prompt
A.6.2 RQ2 Prompts
Generalized Block Execution Prediction Prompt Take a look at the {lang} code blocks. One statement is highlighted in each. You’ll also see the input values going into the function. Based on those, try to figure out what the highlighted line will do. Here are {shot} examples that show how it works: ----------------------------------------
@@ A.6.2 RQ2 Prompts
Generalized input_output prompt Here’s some {lang} code. You’ll either get the inputs or the outputs, but not both. Your task is to fill in the missing part—predict the output if you know the input, or figure out what input must’ve produced the output. Check out these {shot} examples for reference: ----------------------------------------
@@
A.6.3 RQ3 Prompts
Generalized Loop Prediction Prompt Let’s explore some loops in {lang}. You’ll get the full loop structure along with the input values used in the code. I’ll ask you questions about how the loop body or post-loop values behave with those inputs. Here’s how it works with {shot} example(s): ----------------------------------------
@@ A.6.3 RQ3 Prompts
I’ll ask you something about what happens inside the loop body.
@@
I’ll ask you something about what happens after the loop body.

### v2

@@ A.5 Detailed Description of Tasks
In this section, we will provide a detailed description of our tasks. A.5.1 provides a comprehensive description of the statement-based evaluation we performed on LLMs. We sampled five types of statements, i.e, Assignment, Arithmetic, Constant, Boolean, and Function Call, and prompted the models about the value after execution of each type of statement given the variable states before executing that statement. For the block prediction task A.5.2, we sampled statements from the start of the code snippet, given the input of the code, we prompted the model to predict the output at the end of the 1st statement, the 2nd statement, and the 3rd statement. For Branch prediction A.5.3, we promoted the model whether a specific branch will be taken or not of a code snippet, given the input of that code snippet.
@@
In the case of A.5.4, we sampled loop statements from the code snippets. For each loop, we first collected the number of iterations of that loop as ground truth, and queried the model about how many times the loop would be iterated. We also collected and prompted the model regarding the variable state inside the loop body after the n-th interaction. We have named this "In-Loop" prediction. Additionally, we sampled variables after the execution of the whole loop and queried the model regarding the variable value after the execution of the loop body ("Post-Loop" prediction). For input/output prediction A.5.5, we used the approach similar to  Gu et al. [2024]. For output prediction, we give the entire code snippet and the input of the code snippet, and vice versa for input prediction.
A.5.1 Statement Prediction Task
@@
def xldate_from_date_tuple(date_tuple, datemode): year, month, day = date_tuple data_list = list(data_tuple) if datemode not in (0, 1): raise XLDateBadDatemode(datemode) if year == 0 and month == 0 and day == 0: return 0.00 c = 100 if not (1900 <= year <= 9999): raise XLDateBadTuple("Invalid_year:%r" % ((year, month, day),)) if not (1 <= month <= 12): raise XLDateBadTuple("Invalid_month:%r" % ((year, month, day),)) if day < 1 \ or (day > _days_in_month[month] and not (day == 29 and month == 2 and _leap(year))): raise XLDateBadTuple("Invalid_day:%r" % ((year, month, day),)) Yp = year + 4716 M = month if M <= 2: Yp = Yp - 1 Mp = M + 9 else: Mp = M - 3 jdn = ifd(1461 * Yp, 4) + ifd(979 * Mp + 16, 32) + \ day - 1364 - ifd(ifd(Yp + 184, 100) * 3, 4) xldays = jdn - _JDN_delta[datemode] if xldays <= 0: raise XLDateBadTuple("Invalid(year,month,day):%r" % ((year, month, day),)) if xldays < 61 and datemode == 0: raise XLDateAmbiguous("Before1900-03-01:%r" % ((year, month, day),)) return float(xldays) xldate_from_date_tuple(date_tuple=(1907, 7, 3), datemode=0) Assignment Prediction What will be the value of the final output of the statement year, month, day = date_tuple given {’date_tuple’: (1907, 7, 3)} after executing the statement? Arithmetic Prediction What will be the value of the final output of the statement Yp = year + 4716 given {’year’: 1907} after executing the statement? Constant Prediction What will be the value of the final output of the statement c = 0 given {’constant’: 0} after executing the statement? Boolean Prediction Will the true branch of the statement if datemode not in (0, 1): be executed given {’datemode’: 0}? Function Call Prediction What will be the value of the final output of the statement data_list = list(data_tuple) given {’date_tuple’: (1907, 7, 3)} after executing the statement??
A.5.2 Block Prediction Task
@@
A.5.3 Branch Task
1.def xldate_from_date_tuple(date_tuple, datemode): 2. 3. year, month, day = date_tuple 4. 5. if datemode not in (0, 1): 6. raise XLDateBadDatemode(datemode) 7. 8. if year == 0 and month == 0 and day == 0: 9. return 0.00 10. 11. if not (1900 <= year <= 9999): 12. raise XLDateBadTuple("Invalid_year:%r" % ((year, month, day),)) 13. if not (1 <= month <= 12): 14. raise XLDateBadTuple("Invalid_month:%r" % ((year, month, day),)) 15. if day < 1 \ 16. or (day > _days_in_month[month] and not(day == 29 and month == 2 and _leap(year))): 17. raise XLDateBadTuple("Invalid_day:%r" % ((year, month, day),)) 18. 19. Yp = year + 4716 20. M = month 21. if M <= 2: 22. Yp = Yp - 1 23. Mp = M + 9 24. else: 25. Mp = M - 3 26. jdn = ifd(1461 * Yp, 4) + ifd(979 * Mp + 16, 32) + \ 27. day - 1364 - ifd(ifd(Yp + 184, 100) * 3, 4) 28. xldays = jdn - _JDN_delta[datemode] 29. if xldays <= 0: 30. raise XLDateBadTuple("Invalid(year,month,day):%r" % ((year, month, day),)) 31. if xldays < 61 and datemode == 0: 32. raise XLDateAmbiguous("Before1900-03-01:%r" % ((year, month, day),)) 33. return float(xldays) 34. 35.xldate_from_date_tuple((1907, 7, 3), 0) Brach Prediction Is line 12, raise XLDateBadTuple("Invalid year: %r" % ((year, month, day),)) executed when xldate_from_date_tuple((1907, 7, 3), 0) is called?
A.5.4 Loop Task
@@
A.5.5 Input-Output Task
@@
A.6 Concrete to Abstract Mapping
For the approximation of code semantics, we prompted the models to reason in abstract values, instead of reasoning about the concrete exact value. Table 6 shows the mapping from concrete value to abstract category, following the prior literature Ding et al. [2023a]. When defining these mappings, we carefully aligned the value ranges for each abstract category with the overall value distribution observed in our benchmark. We evaluated the abstract mapping results against a random baseline, where mapping rules were selected randomly from all available mapping categories.
@@ A.6 Concrete to Abstract Mapping
$10\leq v<0$ Negative Regular
v<10v<-10 Negative Large
@@
$1.0\leq v<0.0$ Negative Small
$10.0\leq v<-1.0$ Negative Regular
v<10.0v<-10.0 Negative Large
@@ A.7.1 RQ1 Prompt
A.7.2 RQ2 Prompts
Generalized Block Execution Prediction Prompt Take a look at the {lang} code blocks. One statement is highlighted in each. You’$l$ also see the input values going into the function. Based on those, try to figure out what the highlighted line will do. Here are {shot} examples that show how it works: ----------------------------------------
@@ A.7.2 RQ2 Prompts
Generalized input_output prompt Here’s some {lang} code. You’$l$ either get the inputs or the outputs, but not both. Your task is to fill in the missing part—predict the output if you know the input, or figure out what input must’ve produced the output. Check out these {shot} examples for reference: ----------------------------------------
@@
A.7.3 RQ3 Prompts
Generalized Loop Prediction Prompt Let’s explore some loops in {lang}. You’$l$ get the full loop structure along with the input values used in the code. I’$l$ ask you questions about how the loop body or post-loop values behave with those inputs. Here’s how it works with {shot} example(s): ----------------------------------------
@@ A.7.3 RQ3 Prompts
I’$l$ ask you something about what happens inside the loop body.
@@
I’$l$ ask you something about what happens after the loop body.
@@ A.8.3 RQ5
Table 7: Model Performance Comparison on Abstract Value Prediction
Model Name Random Baseline 0-shot 3-shot
DeepSeek-Coder-V2 0.084 0.247 0.504
DeepSeek-R1-Distill-Llama 0.112 0.255 0.549
DeepSeek-R1-Distill-Qwen-14B 0.169 0.316 0.784
DeepSeek-R1-Distill-Qwen-7B 0.113 0.238 0.614
Llama-3.1-8B 0.104 0.266 0.597
Phi-3.5-mini 0.077 0.121 0.577
Phi-4-mini 0.088 0.156 0.478
Qwen2.5-14B 0.176 0.218 0.618
Qwen2.5-Coder-7B 0.099 0.222 0.509
granite-3.2-8b 0.152 0.396 0.528
granite-3.2-8b 0.147 0.367 0.535
A.9 API Definition Ablation Study
In Task 2, when predicting the output for output of an API call, we conducted additional experiments to evaluate whether providing API definitions improves model performance. We tested two types of settings: (i) with No API definition and (ii) with API definitions.
We ran our evaluation on the open-source models and evaluated all the 248 function call prediction examples from our dataset. Table 8 shows the results on the best-performing open-source models:
Table 8: Accuracy of API prediction with different API definition strategies
Model No API definitions API implementation
Qwen 2.5-7B 0.206 0.226
Qwen 2.5-14B 0.182 0.194
Phi-4 0.125 0.089
Llama-3.1-8B 0.105 0.089
The results indicate that providing API definitions does not significantly improve performance, and in some cases slightly degrades it. This supports our hypothesis that the fundamental limitation for fine-grained code reasoning is not lack of API knowledge, but rather the models’ inability to reason about statement-level and block-level semantics.